# Sentient Agent as a Judge: Evaluating Higher-Order Social Cognition in Large Language Models

### v2

@@ 2.2 Sentient Agent as a Judge: Framing Dynamic Environments for Agent Evaluation
When generating each persona $p$, we uniformly sample three characteristics keywords, three sentences that the persona says when chatting with friends, and one persona’s age. We require the base LLM to generate a persona profile based on the given seed information by filling the following slots:
  • Basic Information: Based on the given information, deduce the persona’s name, age, and gender.
  • Occupation, Habits and Daily Behavior: Based on the persona’s information, deduce the persona’s possible occupation and further infer their habits and daily behaviors. Ensure consistency with the persona’s characteristics.
  • Personal Hobbies: Deduce the persona’s personal hobbies, and provide three detailed descriptions that align with the persona’s traits.
  • Speaking Style: Based on the given information and the generated traits, deduce the potential speaking style that matches the persona’s way of communication.
This process ensures the diversity of generated personas for any given task. Additionally, the general seed pools can be replaced with task-specific seed pools when necessary.
@@ 3 Effectiveness of Sage
We demonstrate the robustness of Sage using four distinct sentient agents: DeepSeek-V3, GPT-4o, Gemini 2.5, and Gemini 2.5-Think. We compute pairwise Spearman’s rank correlation coefficients between the rankings for the ten target models listed in Table 1. Rankings generated by different sentient agents exhibit consistently high correlations ($\geq 0.84$), clearly validating the robustness and reliability of Sage. Notably, GPT-4o demonstrates slightly lower alignment with DeepSeek-V3 and Gemini 2.5, whereas the remaining three agents show exceptionally high correlations ($\geq 0.92$). These results reinforce the confidence in and generalizability of Sage, particularly anchored on the open-source DeepSeek-V3.
@@ 4.2 Analysis
By jointly analyzing quantitative economy and qualitative style for the top-10 Arena models, we show today’s strongest models produce empathetic, high-quality responses using significantly fewer tokens. High Sentient scores usually align with an empathy-focused conversational style. These combined insights offer a clearer understanding of LLM social competence, highlighting practical factors (cost, latency) and subtle behavioral traits that single metrics cannot capture.
@@
We use a two-dimensional coordinate system (Figure 7) to evaluate the style of social interaction exhibited by LLMs, complementing the quantitative Sentient score by positioning models based on their orientation (Empathy vs. Solution) and interaction style (Structured vs. Creative). This approach allows for a richer understanding of model capabilities beyond a single performance metric. Based on their performance in supportive dialogues, models are mapped into this 2D space, revealing distinct profiles in how they engage with the user’s emotional state and problems. Construction details are available in Appendix C.
@@ 5 Related Work
The LLM-as-a-Judge paradigm has shown wide applicability across various tasks, including evaluating model performance(Zheng et al., 2023; Qin et al., 2023; Liu et al., 2024; Dubois et al., 2023; Chiang et al., 2024; Wang et al., 2023d; Qin et al., 2024; Zhu et al., 2023; Tu et al., 2024; Tian et al., 2023; Wu et al., 2024; Zhou et al., 2025b), automating data annotation (Taori et al., 2023; Xu et al., 2023; Mukherjee et al., 2023), and providing reward signals(Ouyang et al., 2022; Lee et al., 2023; Chen et al., 2024b; Lightman et al., 2023; Wang et al., 2023a; Hosseini et al., 2024; Snell et al., 2025; Xi et al., 2024; Li et al., 2025).
@@ Appendix B Robustness of Sage
Appendix C Social Cognition Coordinate
  • Model Profile: we first collect ten success cases and ten failure cases for each model, and prompt Gemini2.5-Pro to summarize the reasons behind each success or failure. Then, we construct a profile for each model based on these cases and the overall analysis of the successes and failures.
  • Model Strategy Distribution: we developed a fine-grained strategy list for the Emotional Support Conversation, based on a coarse-grained version from Liu et al. (2021). Next, for each model, we prompt DeepSeek-V3 to analyze each response in all dialogue turns, annotating the strategies used in each response. As a result, we obtain the distribution of strategy usage across all evaluated turns for each model.
Model Profile
Model Strategy Distribution
@@ Appendix C Social Cognition Coordinate
To analyze the support strategies used by an LLM, we prompted DeepSeek-V3 to act as a judge, evaluating each round of the model output to identify all support strategies involved. We then aggregated the strategy statistics across all rounds of conversations, outputing the proportion of rounds each strategy is used. The resulting statistics are then used as one contributing factor in our analysis of the Social Cognition Coordinate system.
@@ Appendix D Case Study
Appendix E Prompt Template for SAGE
@@
Constructing diverse personalities contributes to enhancing the robustness of the benchmark. Therefore, we divide the agent’s profile into two components: persona and background. We meticulously design the attributes that need to be generated to ensure the diversity of profiles.
@@ 5. Use a positive or negative value to represent the character’s emotional changes
Appendix F Prompt Template for External Judger
@@ II. Perceived Attentiveness & Presence
Appendix G Prompt Template for Building Social Cognition Coordinate

### v3

@@ 2.2 Sentient Agent as a Judge: Framing Dynamic Environments for Agent Evaluation
When generating each persona $p$, we uniformly sample three characteristics keywords, three sentences that the persona says when chatting with friends, and one persona’s age. We require the base LLM to generate a persona profile based on the given seed information by filling the following slots: (1) Basic Information: Based on the given information, deduce the persona’s name, age, and gender. (2) Occupation, Habits and Daily Behavior: Based on the persona’s information, deduce the persona’s possible occupation and further infer their habits and daily behaviors. Ensure consistency with the persona’s characteristics. (3) Personal Hobbies: Deduce the persona’s personal hobbies, and provide three detailed descriptions that align with the persona’s traits. (4) Speaking Style: Based on the given information and the generated traits, deduce the potential speaking style that matches the persona’s way of communication.
This process ensures the diversity of generated personas for any given task. Additionally, the seed pools can be replaced with task-specific seed pools when necessary.
@@ 3 Effectiveness of Sage
We validate Sage by demonstrating that the Sentient Agent’s emotions strongly correlate with both internal user thoughts and dialogue quality. Our findings indicate that these emotional scores effectively capture deeper cognitive processes and relational dynamics inherent in supportive interactions.
Setting
We evaluate eight representative LLMs from four major families. For each family, we include both a vanilla model and its corresponding reasoning variant to ensure a balanced and informative comparison:
  • OpenAI: GPT‑4o‑2024‑08‑06 (GPT‑4o, vanilla) and o1‑2024‑12‑27 (OpenAI‑o1, reasoning).
  • DeepSeek: DeepSeek‑V3‑2024‑12‑27 (vanilla) and DeepSeek‑R1 (reasoning).
  • Claude: Claude3.7‑Sonnet, a hybrid model with a toggleable reasoning module. We treat its reasoning-off mode as vanilla, and reasoning-on mode as reasoning.
  • Gemini: Gemini2.5‑Flash, a cost-efficient model that supports both reasoning and non-reasoning modes.
In addition, we include two smaller-scale instruction-tuned open-source models in our analysis: Llama3.3‑70B‑Instruct and Qwen2.5‑72B‑Instruct.
We construct $100$ supportive dialogue scenarios covering $8$ diverse topics to comprehensively evaluate the higher-order social-cognitive abilities of representative LLMs. Unless otherwise specified, we use DeepSeek-V3 as the default sentient agent. Please refer to Appendix A for the details of experimental setting (including BLRI and utterance-level empathy metrics used in the subsequent experiments).
@@
We demonstrate the robustness of Sage using four distinct sentient agents: DeepSeek-V3, GPT-4o, Gemini 2.5, and Gemini 2.5-Think. We compute pairwise Spearman’s rank correlation coefficients between the rankings for the ten target models listed in Table 1. Rankings generated by different sentient agents exhibit consistently high correlations ($\geq 0.84$), clearly validating the robustness and reliability of Sage. Notably, GPT-4o demonstrates slightly lower alignment with DeepSeek-V3 and Gemini 2.5, whereas the remaining three agents show exceptionally high correlations ($\geq 0.92$). These results reinforce the confidence in and generalizability of Sage, particularly anchored on the open-source DeepSeek-V3. Full results can be found in Appendix §B.
Table 2: Human Evaluation Results.
Metrics Scores (%)
Avg. model-human consistency rate 85.3
Model-majority consistency rate 90.0
Avg. inter-annotator agreement 78.5
Avg. reasonableness rating 89.5
Avg. inter-annotator agreement in reasonableness 83.3
Human Evaluation for Sage
We further conducted a human evaluation study to validate the effectiveness of Sage. Specifically, we randomly sampled 100 dialogue contexts and asked 4 human annotators to answer two questions: (1) Based on the given information (identical to the input provided to Sage), what emotional change do you believe the user should experience—an increase or a decrease in emotional score? (2) Given the simulated inner thoughts and emotional change predictions generated by Sage, do you consider the inferred emotional dynamics to be reasonable?
Based on the responses to Question (1), we calculate the average model-human consistency rate, the model-majority consistency rate (i.e., agreement between the model’s predictions and the majority opinion of annotators), and the average inter-annotator agreement. For Question (2), we compute the average reasonableness rating, which reflects how frequently annotators judged the model’s predicted emotional dynamics to be reasonable, along with the corresponding inter-annotator agreement.
As shown in Table 2, Sage achieves high average consistency with human simulations ($85.3\%$), as well as high average reasonableness rating ($89.5\%$), which validates the effectiveness of Sage.
@@ 4.2 Analysis
By jointly analyzing quantitative economy and qualitative style for the top-10 Arena models, we show today’s strongest models produce empathetic, high-quality responses using significantly fewer tokens. High Sentient scores usually align with an empathy-focused conversational style. These combined insights offer a clearer understanding of LLM social competence, highlighting practical factors (cost, latency) and subtle behavioral traits that single metrics cannot capture. We provide more in-depth analysis grounded in emotion score and dialogue utterance including model strategy analysis (§C), case study (§D), and model profile analysis (§E) in Appendix.
@@
We use a two-dimensional coordinate system (Figure 7) to evaluate the style of social interaction exhibited by LLMs, complementing the quantitative Sentient score by positioning models based on their orientation (Empathy vs. Solution) and interaction style (Structured vs. Creative). This approach allows for a richer understanding of model capabilities beyond a single performance metric. Based on their performance in supportive dialogues, models are mapped into this 2D space, revealing distinct profiles in how they engage with the user’s emotional state and problems. Construction details are available in Appendix F.
@@ 5 Related Work
The LLM-as-a-Judge paradigm has shown wide applicability across various tasks, including evaluating model performance (Zheng et al., 2023; Qin et al., 2023; Liu et al., 2024; Dubois et al., 2023; Wang et al., 2023d; Qin et al., 2024; Tu et al., 2024; Tian et al., 2023; Wu et al., 2024; Zhou et al., 2025b), automating data annotation (Taori et al., 2023; Xu et al., 2023; Mukherjee et al., 2023; Chen et al., 2025), and providing reward signals (Ouyang et al., 2022; Lee et al., 2023; Chen et al., 2024b; Lightman et al., 2023; Wang et al., 2023a; Hosseini et al., 2024; Snell et al., 2025; Xi et al., 2024; Li et al., 2025).
@@ Appendix B Robustness of Sage
Appendix C Model Strategy Analysis
C.1 Identifying Model Strategies
@@
To analyze the support strategies used by an LLM, we prompted DeepSeek-V3 to act as a judge, evaluating each round of the model output to identify all support strategies involved. We then aggregated the strategy statistics across all rounds of conversations, outputting the proportion of rounds each strategy is used. The prompt we used for identifying model strategies is presented at §I.
Table 5: Details of the support strategy categorization.
@@
C.2 Strategic Flexibility and Efficiency
Among all evaluated factors, strategic flexibility and efficiency emerges as a key capability for success on the Sentient Leaderboard. Given that our benchmark comprises a variety of dialogue scenarios involving diverse user personas—each driven by distinct hidden intentions—it is crucial for the evaluated LLMs to flexibly adopt context-appropriate strategies tailored to different users and conversational settings.
In this section, we evaluate models’ strategic flexibility and efficiency from two vital perspectives:
  • In-context Strategic Flexibility: The models’ ability to dynamically adjust its strategy within a single dialogue context based on user feedback, rather than relying on repetitive or similar strategy patterns.
  • Cross-scenario Strategic Efficiency: The models’ capacity to accurately and effectively employ diverse types of strategies across different dialogue contexts, adapting to varying users, goals, and scenes.
Figure 9: Strategy Sequence Repetition Rate of Different Models. “(Rank $k$)” represents the rank on the Sentient Leaderboard.
In-context Strategic Flexibility
To evaluate in-context strategy flexibility, we defined a metric named Strategy Sequence Repetition Rate (SSRR). Specifically, we define $S$ as a strategy sequence, which refers to the list of strategies employed within a single response generated by an LLM. For example, the response ”I fully understand your anxiety. Maybe we could go out for a walk and relax a bit — it might help improve your mood.” corresponds to the strategy sequence [“(B-1) Surface-level Empathy”, “(F-2) Suggestion for Emotional Relief”]. Ideally, a flexible model is able to adapt its strategy sequence dynamically in response to user feedback, rather than rigidly adhering to a pre-defined emotional support pattern (which is an approach often associated with less-capable AI systems and their stereotypical behavior). Thus, we define the Strategy Sequence Repetition Rate (SSRR) of each model as follows:
$\text{SSRR}=\frac{1}{|D|}\sum_{d\in D}\left(\frac{1}{N_{d}-1}\sum_{i=1}^{N_{d}-1}\text{G}_{\text{sim}}(S^{d}_{i},S^{d}_{i+1})\right)$
where $D$ is the set of all dialogues of the model and $S^{d}_{i}$ is the strategy sequence of the $i^{th}$ response in dialogue $d$. $\text{G}_{\text{sim}}$ is a similarity measure. Here, we use the Needleman-Wunsch Algorithm (needleman1970general) to obtain the global similarity of the adjacent strategy sequences.
Figure 9 presents the SSRR evaluation results across different models. By setting $\text{SSRR}<0.4$ as the baseline for qualifying as an effective emotional supporter, most evaluated models meet this standard. Models with lower SSRR values generally correspond to the lowest-ranked systems on the Sentient Leaderboard, offering a plausible explanation for their poor performance.
Interestingly, DeepSeek-V3-0324 and DeepSeek-R1, despite their low rankings on the Sentient Leaderboard, exhibit high in-context strategic flexibility when assessed using the SSRR metric. This result partially aligns with our observations in the Social Cognition Coordinate analysis (§F) (it is worth noting that the “structured-to-creative” dimension in the Social Cognition Coordinate incorporates broader considerations beyond in-context strategy flexibility alone) and Case Study findings (§D).
Figure 10: Results of Strategy Efficiency.
Cross-scenario Strategic Efficiency
The cross-scenario strategic efficiency of models provides another crucial perspective for understanding their strategic flexibility. Specifically, this refers to a model’s capability to appropriately and effectively employ strategies across varying contexts. “Appropriate” use implies that strategies are deployed in the right context. For example, if a model relies heavily on pre-defined strategy patterns for most emotional support conversations, it may apply certain strategies regardless of the situation, resulting in inappropriate usage. “Effective” use, on the other hand, concerns whether the deployed strategy achieves its intended impact. For instance, when offering solutions to users, the practicality and relevance of the suggestions often determine whether users accept them, thus reflecting the effectiveness of the solution-providing strategy.
To evaluate the appropriateness and effectiveness of strategy usage, we define the Strategy Effectiveness (SE) of each strategy type as follows:
$\text{SE}=\frac{1}{N}\sum_{i=1}^{N}\ \text{EmoChange}(s_{i})$
where $s_{i}$ denotes an instance of the given strategy type, and $N$ is the total number of such instances within the evaluation context. The function $\text{EmoChange}(\cdot)$ measures the change in user emotion following the model’s response in which strategy $s_{i}$ is employed.
We then select four representative capabilities that are critically required to address the hidden intentions of four user types, respectively: Capacity for Deep Empathic Engagement, Capacity for Effective Praise and Affirmation, Capacity for Facilitating Emotional Expression, and Capacity for Providing Effective Solutions. Each capability corresponds to specific types of important strategies, as defined in Table 5. For instance, the Capacity for Providing Effective Solutions involves the use of the strategy type “(F-5) Advice Specific to the Client’s Situation”. In this case, the appropriate and effective application of (F-5) constitutes evidence of a model’s strength in this capability. The detailed correspondence among capabilities, hidden user intentions, and associated strategy types is presented in Table 6.
@@ C.2 Strategic Flexibility and Efficiency
Capability Related Strategies Related Hidden Intention (task)
Deep Empathic Engagement B-2 You hope the other person will deeply empathize with your feelings, rather than simply offering comfort.
B-3
C-1
C-2
Effective Praise and Affirmation E-1 You hope the other person will sincerely praise your specific actions in the situation
E-2
E-3
Effective Praise and Affirmation D-2 You want the other person to attentively listen to your emotional outpouring
D-3
Providing Effective Solutions F-5 You want to receive advice that can truly help you solve your current difficulties
In Figure 10, we present the average Strategy Effectiveness (SE) results for each capability, evaluated both across all conversations (Overall SE) and within conversations specific to the corresponding tasks (Task-specific SE). Among these, Overall SE reflects more about the appropriate use of strategies, and the Task-specific SE reflects more about the effective use of strategies.
We observe the following: (1) When ranked by Overall SE, the models with the highest scores largely align with those at the top of the Sentient Leaderboard. Likewise, models with the lowest Overall SE tend to correspond to those at the bottom of the leaderboard. This indicates a strong correlation between a model’s success in emotional support conversations and its cross-scenario strategic efficiency, especially the appropriate use of strategies. (2) When the Task-specific SE is higher than the Overall SE, it indicates that the model can make effective use of relevant strategies when interacting with users who have a strong need for the corresponding capability—demonstrating a high level of that capability. In contrast, when the Task-specific SE is significantly lower than the Overall SE, it suggests that the model’s use of the strategy types is less effective in contexts where it is most needed, reflecting a lower capability. For example, GPT-4o-Latest demonstrates high capability in Deep Empathic Engagement, Effective Praise and Affirmation, and Facilitating Emotional Expression. Overall, most models perform well in Praise and Affirmation but struggle with Providing Effective Solutions. Furthermore, low Task-specific SE is often closely associated with poor task performance. For instance, GPT-4.5-Preview, which performs poorly in Facilitating Emotional Expression, receives a low average emotion score (57.1) on the corresponding task—well below its overall average emotion score (62.7). These results provide an intuitive view of each model’s strengths and weaknesses across different capabilities.
@@ Appendix D Case Study
Appendix E Model Profiling
Appendix F Social Cognition Coordinate
  • Model ProfileE): we first collect ten success cases and ten failure cases for each model, and prompt Gemini2.5-Pro to summarize the reasons behind each success or failure. Then, we construct a profile for each model based on these cases and the overall analysis of the successes and failures.
  • Model Strategy DistributionC): we developed a fine-grained strategy list for the Emotional Support Conversation, based on a coarse-grained version from Liu et al. (2021). Next, for each model, we prompt DeepSeek-V3 to analyze each response in all dialogue turns, annotating the strategies used in each response. As a result, we obtain the distribution of strategy usage across all evaluated turns for each model.
Appendix G Prompt Template for Sage
@@
Constructing diverse personalities contributes to enhancing the robustness of the benchmark. We build the agents’ diverse profiles with two components: persona and background. We meticulously design the attributes that need to be generated to ensure the diversity of profiles.
@@ 5. Use a positive or negative value to represent the character’s emotional changes
Appendix H Prompt Template for Analysis
@@ II. Perceived Attentiveness & Presence
Appendix I Prompt Template for Building Social Cognition Coordinate